Suppose that the sound level of a conversation is initially at an angry 71 dB and then drops to a soothing 53 dB. Assuming that the frequency of the sound is 506 Hz, determine the (a) initial and (b) final sound intensities and the (c) initial and (d) final sound wave amplitudes. Assume the speed of sound is 347 m/s and the air density is 1.21 kg/m3.
Answer #1

The intensity level is given by

β = 10*log(I/I0)

I is the intensity of the sound produced

I0 is the threshold of hearing = 10-6 μW/m2

Given initial level of sound βi = 71 dB

final sound level βf = 53 dB

71 = 10*log(Ii/10-6)

initial intensity of the sound Ii = 12.589 μW/m2

53 = 10*log(If/10-6)

If =0.19953  μW/m2

The intensity of the sound is related to the amplitude by the equation

I = 0.5*ρvω2S2

For the initial sound level

12.589*10^-6 = 0.5*1.21*347*(2*pi*506)^2*Si2

Si = 77.023 nm   

For the final sound level

0.19953*10^-6 = 0.5*1.21*347*(2*pi*506)^2*Sf2

sf = 9.69686 nm

****************************************************************************
Goodluck Comment in case any doubt, will reply for sure..